OpenStudy (anonymous):

Bill has a fifty billion dollar. you have a dollar. your money become 2 times larger than a year before. (ex: 1->2->4->8->...) How long year does it take that yours is more than Bill's. give me an answer without calculator.

OpenStudy (anonymous):

I will keep track of my places here so i dont lose place.

OpenStudy (anonymous):

128 7 Years

OpenStudy (anonymous):

1024 10 years

OpenStudy (anonymous):

4096 12 Years

OpenStudy (anonymous):

8192 13 Years

OpenStudy (anonymous):

16384 14 Years

OpenStudy (anonymous):

Sorry if its taking me time im switching between two questions and my work

OpenStudy (johnweldon1993):

Can't you just do \(\large 2^x \ge 50000000000\) ? since x = 0 would be 1 and x = 1 would be 2 etc...

OpenStudy (anonymous):

32768 15 Years

OpenStudy (anonymous):

Yes but if you dont do it in order and just put a inequality then it wont work because a inequality can be any number that would work

OpenStudy (anonymous):

And it would be \[2n > 50000000000\]

OpenStudy (anonymous):

It doesnt say equal to or greater than it says greater than

OpenStudy (anonymous):

Now let me finish my math please

OpenStudy (johnweldon1993):

Right^ but adding the "equal to" will allow us to know when to stop...meaning when we see that x = 35.54 we can obviously say it will take 36 years to be more than Bill's Because as you said...just Greater than..could be any number

OpenStudy (anonymous):

John you have successfully ruined my day. I had one goal and now its ruined.

OpenStudy (anonymous):

2^n=5000000000 log2^n=log5*10^9 nlog2=log5+9log10 n=(log5+9)/log2 so n should be greater than (log5+9)/log2, 32.22years

OpenStudy (anonymous):

Or you could do it my way...

OpenStudy (anonymous):

@eazykang WE'RE TALKING ABOUT YEARS NOT MONTHS SO IT'S MUST BE A NATURAL NUMBER

OpenStudy (anonymous):

Dont yell man its just a number

OpenStudy (anonymous):

in that case u can write 33 years

OpenStudy (anonymous):

btw did question provide log 5 and log 2? or else i think it cannot be solved without using calculator

OpenStudy (anonymous):

Why not?

OpenStudy (anonymous):

I was doing it without calcutlator until i was ruined

OpenStudy (anonymous):

Now if you dont mine i will continue

OpenStudy (anonymous):

65536 16 Years

OpenStudy (anonymous):

131072 17 Years

OpenStudy (anonymous):

262144 19 Years

OpenStudy (anonymous):

524288

OpenStudy (anonymous):

20 Years

OpenStudy (anonymous):

1048576 21 Years

OpenStudy (anonymous):

209722 22 Years

OpenStudy (anonymous):

Thanks for the medal sorry i didnt finish it i had lunch

OpenStudy (anonymous):

419444 23 Years

OpenStudy (anonymous):

828888 24 Years

OpenStudy (anonymous):

838888 sorry typo

OpenStudy (anonymous):

165776 25 Years